# GUS-Net: Social Bias Classification in Text with Generalizations, Unfairness, and Stereotypes

### v2

@@ 4.1 Settings
Given the imbalanced class distribution in our dataset, we evaluated both the macro-average performance of the model and individual entity-type-level metrics. By treating B- and I- tags as a single entity (e.g., combining B-GEN and I-GEN predictions), we enhance our evaluation of the model’s ability to detect the presence of each biased entity, rather than merely assessing the boundaries. This approach allows us to gain deeper insights into the model’s performance across the diverse classes of bias present in the data.
4.1.1 Hardware and Environment
All experiments were conducted on a single NVIDIA T4 GPU with 16GB of memory, hosted on Google Colab, utilizing under 10GB of RAM. The codebase was implemented using PyTorch and the Transformers library, and executed on Ubuntu 20.04 with Python 3.8. We employed pytorch-lightning to streamline the training loops and logging mechanisms.
Metrics | Macro | Entity-type-based | ||||
GUS-Net | Hamming Loss | 0.0528 | Generalizations | Unfairness | Stereotypes | Neutral |
F1 | 0.8 | 0.74 | 0.61 | 0.90 | 0.95 | |
@@ 4.1.1 Hardware and Environment
Recall | 0.77 | 0.72 | 0.49 | 0.90 | 0.97 | |
Nbias | Hamming Loss | 0.06 | Generalizations | Unfairness | Stereotypes | Neutral |
@@
Precision | 0.93 | 0.87 | 0.83 | 0.94 | 0.95 | |
Recall | 0.63 | 0.56 | 0.11 | 0.86 | 0.96 |
4.1.2 Hyperparameters
We trained our BERT-based multi-label token classification model with seven output classes over 17 epochs. The training process utilized a batch size of 16 and an initial learning rate of . The AdamW optimizer with weight decay was implemented, along with a linear learning rate scheduler featuring a warm-up ratio of 0.1. To handle class imbalance, we employed focal loss with for the I-GEN label. The classification threshold for all labels was set at 0.5. The original dataset was partitioned into training (70%), validation (15%), and test (15%) splits, ensuring similar distributions of biased entity types across these splits.
4.2 Results
4.2.1 Overall Performance on Multi-Label Classification for Token-level Bias
The primary goal of this experiment is to evaluate the performance of our model in accurately identifying biased entities at the token-level within the GUS dataset. By focusing on token-level classification, we aimed to capture occurrences of social bias at the level of individual words and phrases, rather than merely at the sentence level.
Due to the absence of directly comparable existing token-level work as a baseline for our model, we opted to implement a variant of Nbias, which was the state-of-the-art (SOTA) method designed for token-level classification. We modified Nbias into a multi-label framework and fine-tuned it on the GUS dataset. This adaptation allowed us to examine how well the Nbias architecture could perform in at multi-label classification, even though it was designed for multi-class classification, while also addressing the challenges associated with imbalanced class distributions. In our implementation, we employed focal loss instead of binary cross-entropy to better manage class imbalance. This choice was critical, as our dataset exhibited significant disparities in the representation of biased entities, particularly with respect to the Unfairness class. From the results in Table 3, we have the following observations:
The Hamming Loss observed for GUS-Net (0.0528) was on par with Nbias (0.564), meaning they classified a similar overall number of labels correctly.
GUS-Net demonstrates better overall F1 (0.80) and Recall (0.77) metrics compared to the F1 (0.68) and Recall (0.63) observed with Nbias. The superior F1 Score and Recall for GUS-Net highlight its effectiveness in identifying presences rather than absences. This could be due to the incorporation of focal loss during training, which allows the model to focus more on difficult-to-classify examples, thereby improving its overall sensitivity to the presence of bias. Conversely, while Nbias achieves higher Precision (0.93) than GUS-Net (0.82), its lower Recall indicates that it may not be capturing all relevant instances of bias.
The detailed F1 scores for each entity type in GUS-Net show strong performance, particularly in the Stereotypes (0.90) and Generalizations (0.74) classes, without sacrificing performance on the Neutral (0.95) class. Our model balances performance across entity types, which suggests the effectiveness of focal loss to encapsulate imbalance of biased entities.
4.2.2 Comparison with Human Annoation
@@
4.2.3 Ablation Study
We conducted an ablation study to evaluate the impact of different configurations on the model’s performance. Table 4 presents the macro-average Precision, Recall, F1-score, and Hamming Loss for the following settings: (1) Our proposed GUS-Net model; (2) GUS-Net without GUS dataset: This configuration relies on an existing corpus, BABE [27]. Since there are no token-level annotations for BABE, we used the same annotation pipeline outlined in this paper. (3) GUS-Net without focal loss: In this configuration, we trained the model using the Binary Cross-Entropy (BCE) loss function.
@@ 4.3 Parameter Sensitivity Study
To identify the optimal focal loss parameters, and , we conducted a sensitivity study by testing various values for each parameter while holding the other constant. As shown in Table 5, we evaluated the performance of the model at different values while keeping fixed at 2. The results indicate that the best-performing value for was , which resulted in improved F1-Scores across all entity types. Table 6 shows the influence of parameter while maintaining at . The results reveal that the macro-average F1-Score remained at , indicating that this combination of parameters effectively balances sensitivity and specificity across entity types. Overall, the sensitivity study highlights the importance of tuning the focal loss parameters to improve the model’s performance in identifying various biases. The optimal values used in this paper ( and ) demonstrate the model’s ability to adapt to class imbalances and enhance its performance in detecting biased entities.
0.1 | 0.2 | 0.4 | 0.65 | 0.8 | |
@@ 4.4 Case Study
To demonstrate our model’s labeling capabilities and generalizability, we present a case study involving religious bias from the GUS dataset. In Figure LABEL:fig:case(a), we provide an example of a statement that exhibits religious bias, along with the corresponding labels generated by GUS-Net. Figure LABEL:fig:case(b) showcases GUS-Net’s outputs for this case study, illustrating its ability to accurately identify and classify instances of religious bias. The outputs are represented visually, highlighting how the model distinguishes between different types of bias, including Generalizations, Unfairness, and Stereotypes. This example indicates the effectiveness of GUS-Net in generalizing across various forms of bias, reinforcing its potential as a robust tool for bias detection in diverse contexts.
5 Conclusion and Discussion

### v3

@@ 4.2 Evaluation Strategy
2.
Hamming Loss: This metric measures the fraction of incorrect labels over all labels in the sequence, accounting for multi-label classification. It is defined as:
where is the total number of tokens multiplied by the number of labels per token, is the true label list for the -th token, is the predicted labels list, and is an indicator function that evaluates whether each true label differs from the prediction labels.
In our evaluation, Hamming loss is effectively similar to accuracy but in the context of multi-label token classification tasks.
@@
Given the imbalanced class distribution in our dataset, we evaluated both the macro-average performance of the model and individual entity-type-level metrics. By treating B- and I- tags as a single entity (e.g., combining B-GEN and I-GEN predictions), we enhance our evaluation of the model’s ability to detect the presence of each biased entity, rather than merely assessing the boundaries. This approach allows us to gain deeper insights into the model’s performance across the diverse classes of bias present in the data, rather than overall accuracy which could show promising metrics even without performing well on our intended entities.
4.3 Evaluation Models and Hyperparameters
Evaluation Strategy
We performed offline evaluation on the test sets using two types of models. The first type includes encoder-only models that were fine-tuned on our specific dataset to adapt to the nuances of our task. The second type comprises decoder-only LLMs, which we explored in two modes: prompting and instruction fine-tuning. We adapted these LLMs to address the challenge of detecting bias at the token-level in textual data. The instruction template for this task would explicitly ask the LLM to identify and classify each token based on its contribution to bias within the text. Below is an example of such an instruction template:
This chat template instructs the LLM to analyze each token in the provided sentence and classify tokens with one or multiple of the GUS entities. The task-specific guidance helps the model to focus on the nuances of language that may convey bias, making it a practical tool for detecting subtle biases in text.
Baseline models Our encoder-only baselines include BERT (trained with focal loss and binary cross-entropy), DistilBERT, and RoBERTa. For decoder-only models, our baselines consist of Llama 3.3, Llama 3.2, Phi 3 Medium, and Phi 3.5 Mini.
We trained our encoder-only multi-label token classification models with seven output classes over 20 epochs. The training process utilized a batch size of 16 and an initial learning rate of . The AdamW optimizer with weight decay was implemented, along with a linear learning rate scheduler featuring a warm-up ratio of 0.1. To handle class imbalance, we used focal loss with and . The classification threshold for all labels was set at 0.5. The original dataset was partitioned into training (75%), validation (15%), and test (10%) splits, ensuring similar distributions of biased entity types across these splits.
For the LLM fine-tuning, a LoRA rank of 16 and of 16 were used, with a dropout of 0. Only Llama 3.3 was quantized to 4-bit, while all other models used 16-bit precision. A temperature of 0.1 was used for the evaluation on a 10% test split of the dataset (the other 90% was used in fine-tuning for 1 epoch). Details of the general hyperparameters and specific QLoRA parameters are provided in Table 3.
| BERT-like Models | LLM Variants |
Models:
BERT-base-uncased, DistilBERT (uncased), RoBERTa (base) |
Models:
Llama3.3-70B, Llama3.2-3B, Phi-3.5-mini-4k, Phi-3-medium-4k |
Learning Rate:
5e-5 with linear scheduler and 10% warm up period |
Learning Rate:
2e-4 with linear scheduler and 5 warm up steps |
Batch Size:
16 |
Batch Size:
8 (Training), 1 (Evaluation) |
Epochs:
20 (early stopping) |
Epochs:
1 |
| Optimizer: AdamW | Optimizer: AdamW |
Weight Decay:
0.01 |
Weight Decay:
0.01 |
Warm-up Steps:
10% of total steps |
Warm-Up Steps: 5 |
Classification Threshold:
0.5 (default for all) |
| LoRA Parameters | |
Parameter: lora_r = 64, lora_alpha = 16, lora_dropout = 0.2
Task Type: CAUSAL_LM Bias: None For Llama 3.3: BitsandBytes: use_4bit = True, bnb_4bit_dtype = float16, bnb_4bit_quant = nf4, use_nested_quant = True |
4.4 Exploratory Data Analysis
We perform an exploratory data analysis on our data and the results are shown in Figure 3. Figures 3(a) and 3(b) depict the distribution of the GUS dataset, in terms of types of bias and entity labels. Figure 3(a) confirms that the dataset is balanced and comprehensive across bias domains. Conversely, Figure 3(b) indicates a class imbalance among entities. Most notably, the neutral entity (O) is present much more often in the dataset than any other token label. Further, some biased entities are much more common than others. For example, stereotypes which are typically multiple tokens long, and are therefore more common in the dataset than unfair tokens, which are usually a single word/descriptor. Entities which are typically longer also contribute to an imbalance between their own B- and I- labels. The domain coverage is balanced, but the inherent label imbalances underline the need for mindfulness during training.
5 Results
In our evaluations, we aim to determine whether encoder-only or decoder-only models are more effective for bias detection in multi-label classification. Our results compare the performance of these model families in recognizing bias for multi-label token classification tasks. The findings are presented and discussed below, with the highest-performing model configuration as GUS-Net: bert-base-uncased, trained with focal loss.
5.1 Comparison between Encoder-only and Decoder-only Models
The analysis presented in the Table 4 focuses on the performance of encoder-only and decoder-only models, specifically fine-tuned on the GUS dataset. This evaluation highlights the effectiveness of each model in bias detection for multi-label classification tasks. Key metrics include Precision, Recall, and F1 score, where higher values indicate better performance. Additionally, the Hamming Loss is reported, with a preference for lower values (below 0.10) to account for class imbalance.
| Model | Hamming Loss | F1 | Precision | Recall | ||
| Encoder-only Models | ||||||
0.05 | 0.80 | 0.82 | 0.77 | ||
| DistilBERT | 0.08 | 0.65 | 0.89 | 0.59 | ||
| RoBERTa | 0.07 | 0.64 | 0.90 | 0.60 | ||
| Nbias (BCE) | 0.06 | 0.68 | 0.93 | 0.63 | ||
| Decoder-only Models | ||||||
| Llama 3.3 | 0.19 | 0.31 | 0.34 | 0.32 | ||
| Llama 3.2 | 0.20 | 0.37 | 0.42 | 0.70 | ||
| Phi 3 Medium | 0.13 | 0.39 | 0.43 | 0.38 | ||
From Table 4, we observe that BERT-base-uncased demonstrates the highest overall performance among encoder-only models, with notable efficiency in balancing precision and recall while maintaining a low Hamming Loss. Conversely, decoder-only models such as Llama 3.2 show higher recall but struggle with overall F1 scores and higher Hamming Loss, indicating potential challenges in handling multi-label classifications. This analysis suggests that encoder-only models may be more suited for this task, especially in contexts where precision in label prediction is critical.
Additionally, Phi 3.5 Mini was evaluated but was unable to generate outputs that could be parsed or aligned with the input. Further observations revealed that while encoder-only models are inherently aligned 100% of the time, the multi-label nature of the task—where each word in the input sequence can have one or many labels—proved challenging for auto-regressive models such as the tested LLMs. For instance, Llama 3.3 output labels were aligned and could be parsed 57.8% of the time, Phi 3 Medium was effective 54.5% of the time, and Llama 3.2 only worked with 14.4% of the inputs, highlighting significant alignment issues with decoder-only models in multi-label token classification tasks.
Key Finding: Encoder-only models, such as BERT model, outperform decoder-only models in terms of both performance metrics and label alignment capabilities, suggesting their superior suitability for complex multi-label classification tasks requiring precise and accurate label predictions.
5.2 Entity-Level Performance of Encoder-only Models
This section evaluates the performance of encoder-only models in identifying biased entities at the token-level within the GUS dataset. We focus on token-level classification to capture occurrences of social bias in individual words and phrases. Table 5 compares encoder-only and decoder-only models.
| Model | Metrics | Macro | Entity-type-based | |||
| Hamming Loss | 0.05 | Generalizations | Unfairness | Stereotypes | Neutral | |
| F1 | 0.80 | 0.74 | 0.61 | 0.90 | 0.95 | |
@@ 5.2 Entity-Level Performance of Encoder-only Models
| BERT-base (GUS-Net) | Recall | 0.77 | 0.72 | 0.49 | 0.90 | 0.97 |
| Hamming Loss | 0.08 | Generalizations | Unfairness | Stereotypes | Neutral | |
| F1 | 0.65 | 0.66 | 0.14 | 0.86 | 0.92 | |
| Precision | 0.89 | 0.87 | 0.85 | 0.94 | 0.90 | |
| DistilBERT | Recall | 0.59 | 0.53 | 0.08 | 0.80 | 0.94 |
| Hamming Loss | 0.07 | Generalizations | Unfairness | Stereotypes | Neutral | |
| F1 | 0.64 | 0.67 | 0.05 | 0.90 | 0.93 | |
| Precision | 0.90 | 0.85 | 0.89 | 0.94 | 0.92 | |
| RoBERTa | Recall | 0.60 | 0.56 | 0.03 | 0.86 | 0.95 |
| Hamming Loss | 0.06 | Generalizations | Unfairness | Stereotypes | Neutral | |
@@
| Precision | 0.93 | 0.87 | 0.83 | 0.84 | 0.93 | |
| Nbias (BCE) | Recall | 0.63 | 0.56 | 0.11 | 0.86 | 0.97 |
Result in Table 5 show that BERT-base (GUS-Net) outperforms other encoder models in almost all metrics, particularly excelling in F1 and Recall, with a notably low Hamming Loss, suggesting high accuracy across multiple label types. This model also shows robust performance across different entity types, even those underrepresented in the dataset, such as Unfairness. In contrast, Nbias, while showing high Precision, falls short in Recall, indicating a potential trade-off between capturing all relevant instances of bias and minimizing incorrect bias predictions. DistilBERT and RoBERTa exhibit lower overall performance, particularly in handling the Unfairness entity.
We summarize our observations as: we observe similar performance between GUS-Net and Nbias, indicating a high level of label classification accuracy. GUS-Net shows superior performance, highlighting its effectiveness in detecting the presence of bias, possibly due to focal loss training that focuses on challenging examples. In terns of entity level performance, GUS-Net demonstrates strong performance across different bias entities, particularly in Stereotypes and Generalizations, without compromising on Neutral entity accuracy. In comparison with decoder-only models, the encoder models, like BERT, show significant advantages in multi-label token classification tasks, providing more detailed and effective bias detection compared to decoder-only models, which struggle with label alignment.
Key Finding Encoder models, particularly when incorporating focal loss, excel in complex multi-label token classification, enabling precise and granular bias detection. These models are not only effective in managing inherent class imbalances but also outperform decoder-only models in alignment and resource efficiency during training and inference.
5.3 Entity-Level Performance of Decoder-only Models
We evaluated decoder-only models using two distinct configurations: prompting and instruction-based fine-tuning (IFT). These approaches were designed to enhance the models’ ability to understand and generate appropriate responses based on specific tasks and instructions, particularly in the context of entity recognition tasks within multi-label classification scenarios. Each configuration aimed to leverage the natural strengths of decoder-only architectures in generating coherent and contextually appropriate
Impact of Instruction Fine-tuning
The evaluation of LLMs instruct fine-tuned on the GUS dataset revealed limitations in their ability to effectively handle multi-label token classification tasks. We performed experiment using Llama3.2 and Llama3.3 and Phi 3 Medium to assess their performance on entity-level.
| Model | Metrics | Macro | Entity-type-based | |||
| Llama 3.3 | Hamming Loss | 0.19 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 216/374 | F1 | 0.31 | 0.21 | 0.11 | 0.40 | 0.50 |
| Precision | 0.34 | 0.24 | 0.13 | 0.30 | 0.71 | |
| Recall | 0.32 | 0.19 | 0.10 | 0.62 | 0.39 | |
| Llama 3.2 | Hamming Loss | 0.20 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 54/374 | F1 | 0.37 | 0.23 | 0.04 | 0.57 | 0.63 |
| Precision | 0.42 | 0.22 | 0.10 | 0.78 | 0.57 | |
| Recall | 0.70 | 0.24 | 0.04 | 0.45 | 0.70 | |
| Phi 3 Medium | Hamming Loss | 0.13 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 204/374 | F1 | 0.39 | 0.25 | 0.01 | 0.58 | 0.72 |
| Precision | 0.43 | 0.30 | 0.06 | 0.67 | 0.69 | |
| Recall | 0.38 | 0.22 | 0.01 | 0.51 | 0.76 | |
Results in Table 6 that Llama 3.3 demonstrated the best alignment ability but exhibited lower accuracy metrics compared to Phi 3 Medium. Phi 3 Medium outperformed other models in terms of F1 Score and Hamming Loss, indicating better overall efficiency in label prediction. Both models struggled with the Unfairness category, particularly in terms of F1 Score and Recall, suggesting a difficulty in accurately identifying less frequent labels. Despite higher alignment, the precision and recall for aligned labels varied significantly, reflecting challenges in consistently predicting correct labels across different entity types. The results highlight the inherent challenges faced by decoder-only models in multi-label token classification, especially concerning label alignment and accurate prediction across diverse entity types.
Key Finding The results highlight the inherent challenges faced by decoder-only models in multi-label token classification, especially concerning label alignment and accurate prediction across diverse entity types.
Impact of Few Shot Prompting
For a complementary approach to instruction-based fine-tuning, we evaluated the Llama 3.3 70B model using few-shot prompting. This method aimed to leverage the model’s ability to generalize from limited examples to manage the task of multi-label token classification. However, alignment of labels with input tokens presented similar challenges as observed with fine-tuning. Notably, when provided with no examples of correct labels, the model failed to align any labels, despite clear formatting instructions in the prompt. This underscores the model’s dependency on example driven guidance for accurate performance.
When employing five and ten-shot prompting, the results, although not as comprehensively aligned as those from fine-tuning, showed improved accuracy on the labels that were correctly aligned. In Table 7, we report a Hamming loss of 0.16 with ten-shot prompting, which reflects a 3% improvement over the fine-tuning approach. Despite this improvement, the performance still does not surpass that of encoder-only models, highlighting the inherent limitations of decoder-only models in handling complex multi-label classification tasks effectively without substantial example guidance.
| Examples in Prompt | Metrics | Macro | Entity-type-based | |||
| 5 shot | Hamming Loss | 0.18 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 169/374 | F1 | 0.37 | 0.21 | 0.18 | 0.38 | 0.70 |
| Precision | 0.38 | 0.19 | 0.17 | 0.50 | 0.66 | |
| Recall | 0.37 | 0.24 | 0.20 | 0.31 | 0.75 | |
| 10 shot | Hamming Loss | 0.16 | Generalizations | Unfairness | Stereotypes | Neutral |
| Aligned: 148/374 | F1 | 0.38 | 0.23 | 0.19 | 0.36 | 0.75 |
| Precision | 0.40 | 0.23 | 0.18 | 0.49 | 0.69 | |
| Recall | 0.38 | 0.23 | 0.20 | 0.28 | 0.81 | |
Qualitative Analysis
Since the auto-regressive nature of LLMs showed limitations in the multi-label token classification task, we prompted them to identify the entities in example inputs without the necessity to classify specific tokens. From the results in Table 8, we can see many LLMs are effective in identifying subtle or implicit bias, though they appear to have a hard time separating the task of understanding a socially biased context and the task of identifying parts of speech such as generalizations, unfairness, or stereotypes.
| Model | Input Text | Entities Predicted and Reasoning | Observation |
| Llama 3.3 | Young men are always causing trouble. |
Entities: Generalization, unfairness, and stereotype.
Rationale: “Overly broad and inaccurate statement perpetuating negative stereotypes about a specific demographic.” |
Correctly identified the explicit bias against young men, aligning with true labels. |
| Llama 3.2 | The women in the office discussed future improvements for the department. |
Entities: None
Rationale: “The statement highlights women in leadership positions, promoting positive representation in the workplace.” |
Missed the opportunity to recognize the context as non-biased. |
| Phi 3 Medium | He’s really articulate for someone from that neighborhood. |
Entities: Generalization, stereotype.
Rationale: “The statement implies a generalization about people from a certain neighborhood’s ability to be articulate.” |
Correctly identified implicit bias despite the positive sentiment. |
| Phi 3.5 Mini | Black programmers created this software. |
Entities: None.
Rationale: Acknowledges diversity in the tech industry. |
Failed to recognize the unnecessary racial marking as bias. |
Due to the superior performance of encoder-only models in our previous setup, we will now be using GUS-Net with BERT setting in subsequent applications.
5.4 Evaluating Model Performance Against Expert-Annotated Bias
@@
5.5 Ablation Study
We conducted an ablation study on GUS-Net to evaluate the impact of different configurations on the model’s performance. Table 9 presents the macro-average Precision, Recall, F1-score, and Hamming Loss for the following settings: (1) Our proposed GUS-Net model; (2) GUS-Net without GUS dataset: This configuration relies on an existing corpus, BABE [18]. Since there are no token-level annotations for BABE, we used the same annotation pipeline outlined in this paper. (3) GUS-Net without focal loss: In this configuration, we trained the model using the binary cross-entropy (BCE) loss function.
@@ 5.6 Parameter Sensitivity Study
To identify the optimal focal loss parameters, , and , we conducted a sensitivity study by testing various values for each parameter while holding the other constant. As shown in Table 10, we evaluated the performance of the model at different values while keeping fixed at 2. The results indicate that the best-performing value for was , which resulted in improved F1 scores across all entity types. Table 11 shows the influence of parameter while maintaining at . The results reveal that the macro-average F1-Score remained at , indicating that this combination of parameters effectively balances sensitivity and specificity across entity types. Overall, the sensitivity study highlights the importance of tuning the focal loss parameters to improve the model’s performance in identifying various biases. The optimal values used in this paper ( and ) demonstrate the model’s ability to adapt to class imbalances and enhance its performance in detecting biased entities.
| 0.1 | 0.2 | 0.4 | 0.65 | 0.8 | |
@@ 5.7 Model Validation Example
To demonstrate our model’s labeling capabilities and generalizability, we present a case involving religious bias from the GUS dataset. In Figure LABEL:fig:case(a), we provide an example of a statement that exhibits religious bias, along with the corresponding labels generated by GUS-Net. Figure LABEL:fig:case(b) showcases GUS-Net’s outputs for this case study, illustrating its ability to accurately identify and classify instances of religious bias. The outputs are represented visually, highlighting how the model distinguishes between different types of bias, including Generalizations, Unfairness, and Stereotypes. This example indicates the effectiveness of GUS-Net in generalizing across various forms of bias, reinforcing its potential as a robust tool for bias detection in diverse contexts.
6 Discussion
6.1 Practical and Theoretical Impact
The GUS-Net framework offers practical implications for various domains, including content moderation, social media analysis, and AI-driven auditing tools. By providing a fine-grained, multi-label token classification approach, GUS-Net enables more precise bias identification, reducing the risk of oversimplification seen in traditional binary classification methods. This work is particularly relevant for regulatory and compliance frameworks, where detailed bias categorization is necessary to ensure transparency and accountability.
GUS-Net’s approach to leveraging generative AI for dataset construction demonstrates a scalable solution for addressing data scarcity in bias detection research. This method can be extended to other sensitive NLP applications, such as fairness assessments in hiring platforms or bias mitigation in automated decision-making systems [27]. The findings also reinforce the importance of employing encoder-based models for bias detection tasks, as they offer more reliable token-level classification [42, 43]. This insight informs future model selection strategies for developers and researchers building AI systems for social bias analysis. In future, the integration of explainability techniques, such as attention visualization and interpretability layers, could further enhance the practical applicability of bias detection models. This would allow policymakers and AI practitioners to better understand model predictions and refine mitigation strategies accordingly.
The theoretical impact of the GUS-Net framework lies in its innovative approach to understanding bias detection in NLP. By redefining the problem as a multi-label token classification task, it offers a more nuanced and granular method for identifying biases in text. This approach allows for the detection of multiple, overlapping biases within a single token or phrase, moving beyond traditional binary or single-label classification methods. As a result, GUS-Net provides a deeper and more comprehensive understanding of bias in language, paving the way for more accurate and context-aware bias detection systems in NLP
6.2 Limitations
Just like any study, we also acknowledge a few limitations. First, the use of synthetic data for bias classification raises questions about how well the model can generalize to real-world scenarios. Second, the dataset has an imbalance in entity labels, with neutral tokens (O-labels) being overrepresented. This could skew the model’s learning process, making it more likely to classify tokens as neutral. Although steps were taken to address this, such as using focal loss, other approaches [44] like weighted sampling or creating more balanced datasets could be explored to further improve results.
Third, while the encoder-only models like BERT performed better that decoder only LLMs for multi-label classification problem, the latter could show stronger reasoning abilities [45]. However, auto-regressive models struggled with aligning predicted labels in multi-label tasks [bender2021danger]. It would be useful to combine the strengths of both model families (encoder-only and decoder-only) for the task [46].
Lastly, while the standard NLP benchmarks are used for evaluation, the ethical challenges of bias detection remains subjective [47]. Human biases might still influence the annotation process, and societal changes could shift how bias is defined. In future, it will be a good practice to focus on making training data more fair and ensuring that annotation frameworks are inclusive and adaptable to evolving perspectives.
7 Conclusion